OpenStudy (anonymous):

Please help! How would you graph this linear system? y= 5/4 x+3 y= 1/4 x-1

OpenStudy (anonymous):

I would only graph the x & y intercepts so for the x-int you make y=0 0=5/4x+3 -3=5/4x -12x=5 x=-5/12 now you have one point (-5/12,0) y-int you make x=0 y=5/4(0) + 3 y=3 now you have the second point (0,3) That is only for the first equation, try do the same urself for the second one.